Ops Meeting Wed 28th September Notes

1. Tracker:
  • Tracker readout seems to be unstable, fine for a bit then not fine for a bit, errors coming in from the VME crate in the readout, not really got to the bottom of it, tried swapping VME crate and another card but problem still recurs. Ray: how do the errors manifest themselves? Ed: DAQ Errors and the run stops; "Error unable to communicate with V977", probably can't talk to the crate and V977 is first on the list. Switched out VME crate and V977; Made another remedy approach at solving the instability issue following diagnosis with Craig, comms bus has been broken by dodgy VLSB. Card is out but no spares, spares coming in with Milorad. Two options: 1. leave it in known unstable state or 2. probably fixed state but with board missing. Decided on (2). Ed has posted to elog the dead channels list. Craig will photograph crate in the hall. Shifters will be authorised to power cycle by pressing button accompanied by appropriate photo. Ed is out Thursday through Sunday. Losing three hundred channels is like losing a plane in station one and five. Note straight tracks only requires two stations in principle and three stations in practice. No On-Call.
  • Note that we need an on-call for DAQ and software.
2. Beamline Water
  • Focus coil cooling moved on to West wall manifolds temporarily. Tomorrow will be moved back permanently and could be done today. Sumitomos in the trench moved onto long lines. MAde changes into the trench to separate heat exchange and FC water. Both sets are up and running so trench and SC magnets have nominal waterflows. Small leak in conventional magnet system, one drip every few minutes. Flow meter installed on the primary side. Pierrick has read back of the temperatures. 18 degrees going in. 2 degrees less going in from the roof but more can be gotten. 17-19 degrees. Would be good to explore possibility to run roof water at 15. May be ongoing issue with heat exchanger on the roof. If we supply water to the trench at 19-20 degrees, should be able to supply water to the magnets at 22 degrees, only couple of degrees temperature change on the heat exchangers. Could be too much glycol in the water - which effects heat capacity. Plan for this evening is to run up the magnets. Start with reference run, then all 200 A run except D1 and D2. TBD. Note risk associated magnet powering in the vault.
  • In principle the beamline water is good, until proven otherwise.
  • Story is - ISIS likely to be tuning up mid-afternoon. May not be beam until later on Thursday. Review on Thursday for a possible 4 pm shift.
  • Plan to do a four o'clock run meeting and consider 4 pm start up. May need to do a beam bump tune. Rogers to liase with Bryan.
  • Search and secure under DC, BLOC to do the beamstop mechanical restraint and emergency stops early afternoon.

Updated by Rogers, Chris over 7 years ago ยท 4 revisions